OpenStudy (anonymous):

Find x if y=-2 4x-6/y=15 A.9/2 B. 5 C.3 D.-6/23

OpenStudy (anonymous):

Are those the only answers? I got 9

OpenStudy (anonymous):

Yea those are the only answers.

OpenStudy (anonymous):

Not sure then, I tried solving twice and I just get 9, sorry.

OpenStudy (anonymous):

No, its okay thanks for trying though.

OpenStudy (anonymous):

Can you show me what math you used to get to 9?

OpenStudy (anonymous):

I know what it is, but could you give me a best guess before I say?

OpenStudy (anonymous):

Oh, he's gone :c

OpenStudy (anonymous):

I did 4x-6/-2=15 Then I multiplied both sides by -2

OpenStudy (anonymous):

I'll just put my work here 4x-6/-2=15 4x+3=15 4x=12 x=3

OpenStudy (anonymous):

I got three, I'm pretty sure it's right... Let me check in a calculator

OpenStudy (anonymous):

Yup, it's three Proof: https://www.wolframalpha.com/input/?i=4x-6%2Fy%3D15if+y%3D-2

OpenStudy (anonymous):

You probably are, idk how the hell I got 9 lmao

OpenStudy (anonymous):

sorry i just saw this and i did it they exact same way as sammii09

OpenStudy (anonymous):

Oh :c

OpenStudy (anonymous):

You are right though. The answer is 3